# On the Enumeration of Generalized Cospectral Mates of Graphs

### v1

@@ 3 Arithmetic Constraints on Matrices in $O_{n}(\mathbb{Q})$
Since $\bar{Q}_{1}$ and $\bar{Q}_{2}$ are integral, the product $\bar{Q}_{1}^{T}\bar{Q}_{2}=\ell^{2}Q_{1}^{T}Q_{2}$ is an integral matrix. It follows from the definition of level that $\ell(Q_{1}^{T}Q_{2})$ divides $\ell^{2}$.For each prime factor $p_{j}$ of $\ell$, the assumptions satisfy the conditions of Lemma 14 with $k=k_{j}$. Thus, $p_{j}\nmid\ell(Q_{1}^{T}Q_{2})$. Since $\ell(Q_{1}^{T}Q_{2})$ divides $\ell^{2}$ but contains no prime factor of $\ell$, it must be that $\ell(Q_{1}^{T}Q_{2})=1$. Therefore, $Q_{1}^{T}Q_{2}$ is an integral orthogonal matrix.
Finally, since $Q_{1},Q_{2}\in O_{n}(\mathbb{Q})$, it follows that $Q_{1}^{T}Q_{2}\in O_{n}(\mathbb{Q})$. An integral orthogonal matrix with row sums equal to 1 must be non-negative, which implies $Q_{1}^{T}Q_{2}$ is a permutation matrix. ∎
@@ 4 Proofs of Theorem 1 and Theorem 2
As observed by Qiu et al. [7], if $H$ is generalized cospectral to $G$ and $Q\in\Gamma(G)$ such that $Q^{T}A(G)Q=A(H)$, then $Q^{T}M(G)e=M(H)e$ and consequently, $Q^{T}\hat{W}(G)=\hat{W}(H)$. The following lemma relates the SNF of $\hat{W}(G)$ to that of the original walk matrix $W(G)$.
@@
Let $G\in\mathcal{F}_{n}$. Then, for every prime factor $p$ of $d_{n}(\hat{W}(G))$, we have $\operatorname{rank}_{p}(\hat{W}(G))=n-1$.
@@
Since $Q_{i}^{T}\hat{W}(G)=\hat{W}(H_{i})$, transposing the equation gives $\hat{W}(G)^{T}Q_{i}=\hat{W}(H_{i})^{T}$. Multiplying by $\ell$, we obtain $\hat{W}(G)^{T}\bar{Q}_{i}=\ell\hat{W}(H_{i})^{T}$. As $\hat{W}(H)$ is an integral matrix and $\ell\equiv 0\pmod{p_{j}^{k_{j}}}$ for each prime $p_{j}$ dividing $\ell$, it follows that $\hat{W}(G)^{T}\bar{Q}_{i}\equiv 0\pmod{p_{j}^{k_{j}}}$. By Lemma 18, we have $\operatorname{rank}_{p_{j}}(\hat{W}(G)^{T})=\operatorname{rank}_{p_{j}}(\hat{W}(G))=n-1$.
@@
The conditions of Lemma 15 are therefore satisfied, implying that $Q_{1}^{T}Q_{2}=P$ for some permutation matrix $P$. It follows that $Q_{2}=Q_{1}P$. By Lemma 4, we conclude that $H_{1}$ is isomorphic to $H_{2}$. ∎
@@
As 2 must be a factor of $d_{n}(W(G))$, let the prime factorization of $d_{n}(W(G))$ be written as $2^{k_{1}}\prod_{j=2}^{m}p_{j}^{k_{j}}$, where $p_{1}=2$ and $p_{2},\dots,p_{m}$ are odd primes. Consider a matrix $Q\in\Gamma(G)$. Since $Q^{T}\hat{W}(G)=\hat{W}(H)$ for some graph $H$ generalized cospectral to $G$, Lemma 8 implies that $\ell(Q)\mid d_{n}(\hat{W}(G))$. By Lemma 17, we have $d_{n}(\hat{W}(G))=\frac{d_{n}(W(G))}{2}$. Consequently, the multiplicity of the factor 2 in $d_{n}(\hat{W}(G))$ is $k_{1}-1$, which implies that $2^{k_{1}}\nmid\ell(Q)$.
@@ 5 An Example and Numerical Evaluation
According to Theorem 2, $G$ can have at most $3$ generalized cospectral mates. By a brute-force search over all graphs of size $10$, we find that $G$ has exactly $3$ generalized cospectral mates, denoted $H_{1},H_{2},$ and $H_{3}$ (see Fig. 1). Thus, $G$ is a tight example attaining the bound given in Theorem 2.
@@
$n$ Controllable $|\mathcal{F}_{n}|$ DGS $\overline{\alpha(G)}$ $\max\alpha(G)$
11 6,705 3,573 2,316 0.7792 19
12 7,708 3,646 2,181 0.9312 29
13 8,441 3,844 2,234 0.9901 39
14 9,061 3,895 2,193 1.0231 32
15 9,427 3,911 2,135 1.112 39
20 9,982 3,979 2,155 1.1586 35
@@
Table 1: Statistics for 10,000 random graphs ($p=0.5$) for each order $n$. The columns report, respectively, the order $n$, the number of controllable graphs, the number graphs in $\mathcal{F}_{n}$, the number of DGS ($\alpha(G)=0$), the average value of $\alpha(G)$ and the maximum value of $\alpha(G)$. The last three columns are computed only for $n\leq 20$ due to the computational complexity of factorizing large numbers.
@@ 6 Conclusion
By leveraging the Smith Normal Form of the walk matrix and the level of rational orthogonal matrices, we proved that distinct levels yield non-isomorphic mates, allowing us to bound the number of mates based on the prime factorization of the last invariant factor of $W(G)$. Our numerical experiments demonstrate that $\mathcal{F}_{n}$ encompasses a substantial proportion of random graphs, approximately $40\%$ as $n$ increases, and we provided a tight example where the derived bound is attained.

### v2

@@ 3 Arithmetic Constraints on Matrices in $O_{n}(\mathbb{Q})$
Since $\bar{Q}_{1}$ and $\bar{Q}_{2}$ are integral, the product $\bar{Q}_{1}^{\mathrm{T}}\bar{Q}_{2}=\ell^{2}Q_{1}^{\mathrm{T}}Q_{2}$ is an integral matrix. It follows from the definition of level that $\ell(Q_{1}^{\mathrm{T}}Q_{2})$ divides $\ell^{2}$.For each prime factor $p_{j}$ of $\ell$, the assumptions satisfy the conditions of Lemma 14 with $k=k_{j}$. Thus, $p_{j}\nmid\ell(Q_{1}^{\mathrm{T}}Q_{2})$. Since $\ell(Q_{1}^{\mathrm{T}}Q_{2})$ divides $\ell^{2}$ but contains no prime factor of $\ell$, it must be that $\ell(Q_{1}^{\mathrm{T}}Q_{2})=1$. Therefore, $Q_{1}^{\mathrm{T}}Q_{2}$ is an integral orthogonal matrix.
Finally, since $Q_{1},Q_{2}\in O_{n}(\mathbb{Q})$, it follows that $Q_{1}^{\mathrm{T}}Q_{2}\in O_{n}(\mathbb{Q})$. An integral orthogonal matrix with row sums equal to 1 must be non-negative, which implies $Q_{1}^{\mathrm{T}}Q_{2}$ is a permutation matrix. ∎
@@ 4 Proofs of Theorem 1 and Theorem 2
As observed by Qiu et al. [7], if $H$ is generalized cospectral to $G$ and $Q\in\Gamma(G)$ such that $Q^{\mathrm{T}}A(G)Q=A(H)$, then $Q^{\mathrm{T}}M(G)e=M(H)e$ and consequently, $Q^{\mathrm{T}}\hat{W}(G)=\hat{W}(H)$. The following lemma relates the SNF of $\hat{W}(G)$ to that of the original walk matrix $W(G)$.
@@
Let $G\in\mathcal{F}_{n}$, then for every prime factor $p$ of $d_{n}(\hat{W}(G))$, we have $\operatorname{rank}_{p}(\hat{W}(G))=n-1$.
@@
Since $Q_{i}^{\mathrm{T}}\hat{W}(G)=\hat{W}(H_{i})$, transposing the equation gives $\hat{W}(G)^{\mathrm{T}}Q_{i}=\hat{W}(H_{i})^{\mathrm{T}}$. Multiplying by $\ell$, we obtain $\hat{W}(G)^{\mathrm{T}}\bar{Q}_{i}=\ell\hat{W}(H_{i})^{\mathrm{T}}$. As $\hat{W}(H)$ is an integral matrix and $\ell\equiv 0\pmod{p_{j}^{k_{j}}}$ for each prime $p_{j}$ dividing $\ell$, it follows that $\hat{W}(G)^{\mathrm{T}}\bar{Q}_{i}\equiv 0\pmod{p_{j}^{k_{j}}}$. By Lemma 18, we have $\operatorname{rank}_{p_{j}}(\hat{W}(G)^{\mathrm{T}})=\operatorname{rank}_{p_{j}}(\hat{W}(G))=n-1$.
@@
The conditions of Lemma 15 are therefore satisfied, implying that $Q_{1}^{\mathrm{T}}Q_{2}=P$ for some permutation matrix $P$. It follows that $Q_{2}=Q_{1}P$. By Lemma 4, we conclude that $H_{1}$ is isomorphic to $H_{2}$. ∎
@@
As 2 must be a factor of $d_{n}(W(G))$, let the prime factorization of $d_{n}(W(G))$ be written as $2^{k_{1}}\prod_{j=2}^{m}p_{j}^{k_{j}}$, where $p_{1}=2$ and $p_{2},\dots,p_{m}$ are odd primes. Consider a matrix $Q\in\Gamma(G)$. Since $Q^{\mathrm{T}}\hat{W}(G)=\hat{W}(H)$ for some graph $H$ generalized cospectral to $G$, Lemma 8 implies that $\ell(Q)\mid d_{n}(\hat{W}(G))$. By Lemma 17, we have $d_{n}(\hat{W}(G))=\frac{d_{n}(W(G))}{2}$. Consequently, the multiplicity of the factor 2 in $d_{n}(\hat{W}(G))$ is $k_{1}-1$, which implies that $2^{k_{1}}\nmid\ell(Q)$.
@@ 5 An Example and Numerical Evaluation
According to Theorem 2, $G$ can have at most $3$ generalized cospectral mates. By a brute-force search over all graphs of size 1010, we find that $G$ has exactly 33 generalized cospectral mates, denoted $H_{1},H_{2},$ and $H_{3}$ (see Fig. 1). Thus, $G$ is a tight example attaining the bound given in Theorem 2.
@@
$n$ Controllable $|\mathcal{F}_{n}|$
11 6,705 3,573
12 7,708 3,646
13 8,441 3,844
14 9,061 3,895
15 9,427 3,911
20 9,982 3,979
@@
Table 1: Statistics for 10,000 random graphs ($p=0.5$) for each order $n$. The columns report, respectively, the order $n$, the number of controllable graphs and the number graphs in $\mathcal{F}_{n}$.
@@ 6 Conclusion
By leveraging the Smith Normal Form of the walk matrix and the level of rational orthogonal matrices, we proved that distinct levels yield non-isomorphic mates, allowing us to bound the number of mates based on the prime factorization of the last invariant factor of $W(G)$. Our numerical experiments demonstrate that $\mathcal{F}_{n}$ encompasses a substantial proportion of random graphs, approximately $39\%$ as $n$ increases, and we provided a tight example where the derived bound is attained.